The length of a rectangular field is twice its breadth. Jamal jogged around it four times and covered a distance of 6km. What is the length of the field?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

From the question it is given that,

The length of a rectangular field is twice its breadth = 2b Perimeter of rectangular field = 6km/4

= 1.5 km
We know that, perimeter of rectangle = 2 × (length + breadth)

1.5 = 2 × (2b + b) 1.5
= 2 × 3b
1.5 = 6b
b = 1.5/6

b = 0.25 km length of rectangle field
= 2 × 0.25
= 0.5km

= 0.5 × 1000
= 500 m


Thanks